# Weakly supervised image segmentation for defect-based grading of fresh produce

### v1

@@ Abstract
Implementing image-based machine learning in agriculture is often limited by scarce data and annotations, making it hard to achieve high-quality model predictions. This study tackles the issue of postharvest quality assessment of bananas in decentralized supply chains. We propose a method to detect and segment surface defects in banana images using panoptic segmentation to quantify defect size and number. Instead of time-consuming pixel-level annotations, we use weak supervision with coarse labels. A dataset of 476 smartphone images of bananas was collected under real-world field conditions and annotated for bruises and scars. Using the Segment Anything Model (SAM), a recently published foundation model for image segmentation, we generated dense annotations from coarse bounding boxes to train a segmentation model, significantly reducing manual effort while achieving a panoptic quality score of 77.6%. This demonstrates SAM’s potential for low-effort, accurate segmentation in agricultural settings with limited data.
keywords:
Machine Learning , Computer Vision , Food Quality , Postharvest
@@ 1 Introduction
Visual inspection of agricultural products plays a crucial role in postharvest supply chains. Traditionally, this process has relied heavily on manual inspection, which is labor-intensive, time-consuming, and prone to human error and inconsistency. The inspectors also need dedicated training, for example, to identify different defects, their size, and their severity. Typical defects for banana fruit are, for example, surface bruises and scars, insect damage, sunburn, chilling injuries, malformed fruits, or maturity stains. As a result, there has been a growing interest in leveraging advanced technologies such as computer vision and machine learning to enhance the accuracy and efficiency of visual inspection and grading of agricultural products (Yuvaraj et al., 2023; Barthwal et al., 2024). That way, inspectors with limited training could still accurately identify different defects and the severity of these defects in the field via a smartphone, which is very scalable, also to remote areas.
Machine learning-based computer vision models can establish an automated assessment procedure that increases the speed and consistency of grading. Traditionally, these AI models, however, require a considerably large and well-annotated training dataset to work well in real-life scenarios. However, large parts of the local, regional, and export postharvest supply chain are decentralized, non-automated, and not digitalized from farm to retail. Therefore, stakeholders often do not have large annotated datasets readily available for their specific use case, defects, and cultivars that could be used to set up these AI models (De-Arteaga et al., 2018). Recent computer vision research shows a trend toward foundation models that are pre-trained on large (often unlabeled) datasets, enabled by advances in self-supervised learning (Balestriero et al., 2023). These models can be more easily adapted to complex downstream tasks by requiring less fine-tuning data compared to the previous generation of models and, therefore, facilitate the implementation of machine learning solutions in low data regimes (Knott et al., 2023). Most foundation models can be adapted to various tasks in a specific data modality (e.g., text, image, or time series) or across modalities (e.g., Li et al., 2024; Zhang et al., 2023). Other foundation models are designed to excel in a certain task or domain problem within a modality (e.g., Israel et al., 2023; Shah et al., 2023; Lu et al., 2024).
In this work, we aim to demonstrate that visual foundational models can be used to detect defects based on a relatively low number of images and with coarse annotations only. We address a common type of visual inspection: recognizing surface defects such as blemishes caused by external impacts (Abbas et al., 2019). Our study focuses on banana fruits and images taken via non-static cameras (e.g., smartphones) in the field and in sorting and packing facilities. We identified two key indicators relevant for grading: the number and size of blemishes visible on a bunch. In addition, the type of the defect is of secondary interest as it can help supply chain operators identify the cause of the defect. Such a use case is typically approached via Image Segmentation, a subcategory of Computer Vision that requires training data with dense image annotations, meaning each pixel has to be associated with a category and/or instance, traditionally leading to a tedious annotation process. To address this, our method utilizes the Segment Anything Model (SAM, Kirillov et al., 2023), a visual foundation model that specializes in image segmentation. We use SAM to generate dense image annotations from coarse high-level annotations (bounding boxes) and use them to train another image segmentation model that identifies, classifies, and segments banana surface defects in an image. We demonstrate that this method works well on a small dataset of 476 coarsely annotated images.
2 Materials and Methods
@@ 2.1 Data collection
We collected 476 images of single banana bunches containing at least one visible bruise or scar in three different sorting and packing facilities in India located in Tembhurni (Maharashtra state), Raver (Maharashtra state), and Kamrej (Gujarat state). The images were taken with different smartphone cameras between harvesting and packing. Due to the small dataset size, these images are split into five subsets used for cross-validation. This means that we evaluate all results on five different train-validation splits to ensure the robustness of our results. To ensure independent data samples, each bunch was only photographed once. The photos have realistic, non-standardized properties: the proximity to the fruits, lighting conditions, image sharpness, and resolution can vary. Also, there can be other bananas and objects occasionally in the background. Such conditions are typical for images taken in supply chains by inspectors, so we wanted our method to accommodate that. We aim to first automatically detect, segment, and classify defects using a machine learning model. In the second step, we derive the number of visible surface defects and their relative size from these predictions.
There are different principal approaches to defining target labels in image segmentation. One can assign each pixel to a general class (“semantic segmentation”) or to both a class and an instance (“instance segmentation”). The choice depends on the use case. This work uses a mixed approach called “panoptic segmentation”, which we will justify in the following paragraphs. See Section 2.2 for a more detailed explanation of the segmentation concepts.
Figure 1: Defect examples. Example images of banana surface defects we aim to detect in this study. Bruises usually result from a dull impact, while scars are caused due to a sharp impact. Old defects are usually darker than new defects as they oxidize over time.
@@
The images were then annotated by a single expert using the open-source tool “LabelStudio” (Tkachenko et al., 2020). As we are interested in the number of defects, we annotated defects on an instance level, meaning that we annotated one bounding box per defect. While a key property of our approach is to only use bounding boxes and generate instance masks via SAM, we also annotate defect masks for validation purposes. Specifically, each defect was one of four categories: “Old Bruise”, “New Bruise”, “Old Scar”, and “New Scar”.
Bruises typically result from blunt force or pressure, which can arise from various factors such as handling, packing, impacts, or weight. Scars, on the other hand, are caused by sharp contact due to reasons like friction from leaves or stems, cuts from knives, crates, rough handling, laborers’ nails, or the sharp tips and crowns of neighboring bunches. Old defects are usually darker than new defects as they oxidize over time (see Figure 1). While the automated detection of generic defects without classification is already an impactful use case, we chose to add this categorization as it helps supply chain managers trace back the cause of the damage in the supply chain. It is important to note that there are cases that cannot be unambiguously assigned to one of those categories, even by expert annotators. However, a sub-goal of this work is to assess to which degree an automated classification is possible. It is worth mentioning that there are plenty of other relevant surface defects, reducing the quality of banana crops that are not included in this study, e.g., insect damage, speckles, sunburn, maturity stains, malformed fruits, broken necks, or chemical residues.
@@
To obtain relative defect sizes, we also annotated segmentation masks of banana fruits. We decided to only annotate bananas semantically on a pixel level, omitting instance information for two reasons: First, as images often contain multiple bunches that are partly cropped or occluded, it is challenging to differentiate instances. Second, our use case does not require us to identify instances, as we are not interested in counting single bananas or bunches. We would rather distinguish focused foreground bunches from those that may appear in the background and, therefore, annotate two types of pixel classes: “Foreground Banana” and “Background Banana”. All pixels not annotated as a banana of defect are assigned the “Background” class. We annotated banana masks using LabelStudio and Segment Anything point prompts (see Section 2.3 for more details).
Figure 2: Overview of our approach. We utilize the Segment Anything Model (SAM), a promptable visual foundation model for image segmentation, to generate dense annotations (pixel-wise class and instance labels) from coarse annotations (bounding boxes and reference points) without any model training involved. These dense annotations otherwise require tedious hand-annotation. Using these newly generated labels, we train a panoptic segmentation model to identify surface defects on banana fruits, specifically bruises and scars. Additionally, the model can differentiate between foreground banana fruits and those in the background. This enables us to determine the number and size of visible surface defects from photographs of banana fruits. We validate our approach using a dataset of 476 images.
@@ 2.2 Image segmentation
Image segmentation uses computer vision techniques to analyze and understand images on a pixel level. In this section, we describe different paradigms of image segmentation and their corresponding evaluation metrics that we use to evaluate the quality of our models.
2.2.1 Semantic Segmentation
“Semantic segmentation” is concerned with classifying each pixel of an image into a predefined category (e.g., background, banana fruit, defect) without distinguishing between individual objects (instances) of the same class. (Ronneberger et al., 2015; Chen et al., 2017; Xie et al., 2021)
Semantic segmentation maps are evaluated against a groundtruth using the Intersection-over-Union (IoU) score for a single class or the mean IoU (mIoU) score for all $N$ classes in a dataset (Everingham et al., 2012; Lin et al., 2014). Eqs. 1 and 2 show the definition of those metrics where $G$ refers to the set of groundtruth pixels that are assigned to a class and $P$ to the set of prediction pixels accordingly.
$IoU(G,P)$ $=\frac{|G\cap P|}{|G\cup P|}$ (1)
$mIoU$ $=\frac{1}{N}\sum_{i=1}^{N}IoU_{i}$ (2)
2.2.2 Instance Segmentation
Instance segmentation goes a step further by categorizing each pixel and identifying each object instance of the same class separately. This is realized by assigning each pixel to one of potentially multiple objects and to a semantic class simultaneously. This technique is essential for tasks requiring counting or differentiating between items. It is considered more complex, combining object detection with pixel classification. (He et al., 2017; Lin et al., 2017)
In terms of evaluation, instance segmentation is equivalent to object detection, which in machine learning describes the localization of instances via rectangular bounding boxes. The only difference is that the overlap of instances is defined by masks rather than boxes. There are several widely recognized metrics that compare a set of groundtruth bounding boxes or instance masks to a set of predicted bounding boxes or instance masks. Generally, a predicted and an annotated instance are matched if their localization exceeds a certain IoU threshold. As a first step, we calculate the precision $P$ and recall $R$ as follows:
$P$ $=\dfrac{TP}{TP+FP}$ (3)
$R$ $=\dfrac{TP}{TP+FN}$ (4)
where $TP$ is the number of true positives, $FP$ is the number of false positives, and $FN$ is the number of false negatives. By definition, precision and recall are inversely related but must concur effectively to ensure that both relevant and correct instances are identified. The primary evaluation metric for object detection is Average Precision ($AP$) for a single class and mean Average Precision ($mAP$) for multiple classes. We follow established practices and report three different variations:
  • $AP@50$: $AP$ at a $IoU$ threshold of $0.50$ as the most common metric in many benchmarks (e.g., Everingham et al., 2012),
  • $AP@75$: $AP$ at a $IoU$ threshold of $0.75$ as a more strict metric,
  • $AP$ (or $AP^{IoU=.50:.05:.95}$): $AP$ averaged over ten equidistant $IoU$ thresholds between $0.50$ and $0.95$ as proposed by Lin et al. (2014).
The average recall is another evaluation metric used to measure the assertiveness of object detectors for a given class (Padilla et al., 2021). Similar to $AP$, the $AR$ metric evaluates a range of IOU thresholds. We use the equation provided by Hosang et al. (2015) to calculate $AR$ from a discrete set of thresholds based on the total number of ground truth instances $G$ in an image:
$AR=\dfrac{2}{G}\sum_{i=1}^{G}max(IoU_{i}-0.5,0)$ (5)
2.2.3 Panoptic Segmentation
In scenarios where one requires both instance-level information for some classes that are countable (widely referred to as “things”) and semantic information about other non-countable components (referred to as “stuff”), these two concepts need to be combined into an approach called “Panoptic Segmentation”. This approach seeks to overcome the limitations of its predecessors by delivering a comprehensive segmentation map that includes both class labels and instance identification. (Cheng et al., 2021, 2022; Kirillov et al., 2019a, b)
In our study, we declare defects as “things” and bananas as “stuff” as we are not interested in distinguishing single banana fruits. Thus, we simplify the machine learning task.
Kirillov et al. (2019b) propose Panoptic Quality ($PQ$) as a comprehensive metric for panoptic image segmentation (Eq. 6). The intuition behind $PQ$ is to provide a unified measure that captures the accuracy of detecting and correctly segmenting individual objects (defects in our case) as well as the overall segmentation quality of the entire scene. $PQ$ combines precision (how many predicted segments are correct) and recall (how many true segments are detected) into a single score, rewarding models that can accurately identify object boundaries and classify them correctly.
$PQ=\frac{\sum_{(p,g)\in TP}\text{IoU}(p,g)}{|TP|+\frac{1}{2}|FP|+\frac{1}{2}|FN|}$ (6)
In this work, we use Panoptic Quality as the primary evaluation metric to compare how well trained image segmentation models perform.
2.3 Segment Anything Model (SAM)
The Segment Anything Model (SAM) is a recently published foundation model for image segmentation (Kirillov et al., 2023). It was trained on a large-scale dataset comprising diverse images from various domains, including natural scenes, urban environments, medical imagery, and other contexts. It is the first model of its kind specifically tailored for generalized image segmentation and promises to facilitate the implementation of domain-specific segmentation models.
Figure 2 (left) shows a high-level overview of SAM’s components. The model consists of three major components: (1) a vision transformer (Dosovitskiy et al., 2020) pre-trained with the Masked Autoencoder method (He et al., 2021) that serves as an image encoder; (2) a prompt encoder that allows the user to specify bounding boxes or points to localize the region of interest in an image [^1] In their paper, the authors of SAM mention open vocabulary text as input prompts. However, they did not publish their code for this feature as the results are not as good as for other prompts (REF needed). Therefore, we do not consider this option in this research.; (3) a mask decoder that generates segmentation masks based on the concatenated image and prompt embeddings. This approach yields impressive results for zero-shot instance segmentation in various domains. However, compared to conventional segmentation models, it comes with two disadvantages that require further adaptations or post-processing to enable SAM for a typical image segmentation use case: Firstly, SAM requires weak labels such as bounding boxes or point coordinates as input prompts. Secondly, the output masks are uncategorized, which hinders zero-shot usage of SAM in a multi-class setting.
SAM2 (Ravi et al., 2024) is a recent extension of SAM to the video domain. Although videos are not the focus of this work, SAM2 has demonstrated that their video-based pre-training enhances zero-shot mask generation for images as well. In our study, we assess both SAM and SAM2 using a ViT-L backbone for zero-shot mask generation of banana fruits and defects.
2.4 Experimental setup
Figure 2 shows an overview of our approach. The process starts with an image dataset with coarsely annotated labels: bounding boxes to locate and assign labels to defects and sets of points that identify foreground and background bananas. We utilize a promptable foundation model for image segmentation (SAM) to generate fine-grained pixel-wise annotations of each image in the dataset based on the coarse hand annotations. It is important to note that SAM is not trained in this procedure but is used here in a zero-shot manner to infer dense annotations only. These newly created annotations are used to train a panoptic segmentation model in a standard supervised machine learning workflow. The panoptic segmentation model we use is a Maskformer (Cheng et al., 2021) model pre-trained on the ADE20k dataset (Zhou et al., 2017). Further details on model training, such as hyperparameters and data augmentation, can be found in A.
@@ 2.4 Experimental setup
Validating SAM masks.
Qualitative analysis indicated that SAM-generated defect masks are generally plausible. However, to quantify the impact of training on SAM-generated masks compared to using hand-annotated pixel-wise masks, we repeat our experiments with different combinations of mask sources for training and validation data: (a) both training and validation use hand-annotated defect masks; (b) both training and validation use SAM-generated defect masks; and (c) training uses SAM-generated masks while validation uses hand-annotated masks. By comparing the evaluation metrics across these three variations, we can quantify the trade-off between the time-saving benefits of automated mask generation and the quality of the resulting output.

### v2

# Weakly Supervised Panoptic Segmentation for Defect-Based Grading of Fresh Produce

@@ Abstract
Visual inspection for defect grading in agricultural supply chains is crucial but traditionally labor-intensive and error-prone. Automated computer vision methods typically require extensively annotated datasets, which are often unavailable in decentralized supply chains. We address this challenge by evaluating the Segment Anything Model (SAM) to generate dense panoptic segmentation masks from sparse annotations. These dense predictions are then used to train a supervised panoptic segmentation model. Focusing on banana surface defects (bruises and scars), we validate our approach using 476 field images annotated with 1440 defects. While SAM-generated masks generally align with human annotations, substantially reducing annotation effort, we explicitly identify failure cases associated with specific defect sizes and shapes. Despite these limitations, our approach offers practical estimates of defect number and relative size from panoptic masks, underscoring the potential and current boundaries of foundation models for defect quantification in low-data agricultural scenarios. GitHub: https://github.com/manuelknott/banana-defect-segmentation
footnotetext: Accepted as a paper to the 6th International Workshop on Agriculture-Vision: Challenges & Opportunities for Computer Vision in Agriculture in conjunction with IEEE/CVF CVPR 2025.
@@ 1 Introduction
Visual inspection of agricultural products plays a crucial role in postharvest supply chains. Particularly, fresh produce is often graded based on visible blemishes and defects. Traditionally, this process relies on manual inspection—an approach that is labor-intensive, time-consuming, and prone to inconsistency and human error. Additionally, inspectors require specialized training to recognize various types of defects, assess their size, and determine their severity. For those reasons, machine learning-based computer vision models have become increasingly popular for automated assessment, increasing the speed and consistency of grading [28, 3]. These deep learning models traditionally require a considerably large and well-annotated training dataset to work well and robustly in real-world scenarios. However, large parts of the local, regional, and export postharvest supply chain are decentralized, non-automated, and not digitalized from farm to retail. Therefore, stakeholders often do not have large annotated datasets readily available for their specific use case, defects, and cultivars that could be used to set up these models [6]. We argue that this domain has a particular need for computer vision models that work in the low-data regime, where images, annotations, or both are limited.
Recent computer vision research shows a trend toward foundation models that are pre-trained on large (often unlabeled) datasets, enabled by advances in self-supervised learning [2]. From a practical view, these models promise to address low-data regimes as they are more easily adaptable to complex downstream tasks in a zero or few-shot manner [16]. Most foundation models can be adapted to various tasks (e.g., classification, detection, segmentation, depth estimation) in a specific data modality (e.g., images, text, time series) or across modalities [e.g., 17, 29]. Other foundation models are designed to excel in a certain task [15, 27] or domain problem [11, 22, 19] within a modality.
In this work, we address the defect-based grading of fresh produce under constrained data and annotation availability, which occurs in many real-world scenarios, particularly in decentralized supply chains. Our end goal is to count and estimate the size of visible surface defects (i.e., surface blemishes such as bruises and scars) on fresh produce, which is essential for grading. We formulate this problem as a panoptic segmentation task [14]. While our experiments focus on bananas as a representative commodity, the proposed methodology is generalizable to other types of fresh produce with comparable grading criteria.
Our contribution is two-fold: First, we address the issue of label scarcity that naturally occurs with dense computer vision tasks such as semantic or panoptic segmentation. We will evaluate the Segment Anything Model (SAM) [15], a promptable foundation model for image segmentation, for the weakly supervised generation of panoptic masks. More specifically, we evaluate how defect instance masks align with dense human annotations and look for systematic failure cases. Second, using these predicted masks, we train a panoptic segmentation model on 476 images. We evaluate results against both human-annotated ground truth and masks generated by SAM. This allows us to estimate how the error introduced by using SAM-generated masks propagates through the downstream task. To our knowledge, this is one of the first works to combine foundation models like SAM with panoptic segmentation for defect quantification in the agricultural domain.
2 Methods
Figure 1: Overview of our approach. We utilize the Segment Anything Model (SAM), a promptable visual foundation model for image segmentation, to generate dense annotations (pixel-wise class and instance labels) from coarse annotations (bounding boxes and reference points) without any model training involved. These dense annotations otherwise require tedious hand-annotation. Using these newly generated labels, we train a panoptic segmentation model to identify surface defects on banana fruits, specifically bruises and scars. Additionally, the model can differentiate between foreground banana fruits and those in the background. This enables us to determine the number and size of visible surface defects from photographs of banana fruits. We validate our approach using a dataset of 476 images and 1440 annotated defects.
2.1 Framing the problem
Our work focuses on a common form of visual inspection relevant across a wide range of commodities: detecting surface defects, such as blemishes resulting from external impacts [1]. Although grading criteria vary by commodity and market, we identify two key indicators that are consistently important: (a) the number of visible blemishes and (b) the size of each of those. Additionally, the type of defect, while secondary, provides valuable insight into potential causes, aiding supply chain traceability and intervention. We argue that an image segmentation approach is better suited for this task than regression or ordinal classification. Segmentation outputs are inherently interpretable and can be visually validated, which is crucial in quality assurance settings where transparency and accuracy are essential.
To estimate the number of defects, it is essential to distinguish between individual defect instances. Additionally, identifying banana fruits is necessary for relative size estimation (see Section 2.6 for details); however, distinguishing between different banana fruit instances is not required beyond distinguishing foreground and background bunches in an image. Based on these requirements, we conclude that the panoptic segmentation [14] framework—combining both semantic and instance segmentation—is well-suited to our task. In principle, a standard semantic segmentation approach combined with postprocessing could also be used to separate instances (see Section 2.5), but panoptic segmentation offers a more integrated and efficient solution.
@@ 2.2 Data collection
We collected 476 images of single banana bunches containing at least one visible bruise or scar in three different sorting and packing facilities in India, located in Tembhurni (Maharashtra state), Raver (Maharashtra state), and Kamrej (Gujarat state). The images were taken with different smartphone cameras between harvesting and packing. To ensure independent data samples, each bunch was photographed only once. The photos have realistic, non-standardized properties: the proximity to the fruits, lighting conditions, image sharpness, and resolution can vary. Also, there can be other bananas and objects occasionally in the background. Such conditions are typical for images taken in supply chains by inspectors, so we wanted our method to accommodate that. We aim to first automatically detect, segment, and classify defects using a machine learning model. In the second step, we derive the number of visible surface defects and their relative size from these predictions.
@@
The images were annotated by a single expert using the open-source tool “LabelStudio” [24]. As we are interested in the number of defects, we annotated defects on an instance level, meaning that we annotated one bounding box per defect. While a key property of our approach is to only use bounding boxes and generate instance masks via SAM, we also annotate defect masks for validation purposes. Specifically, each defect was one of four categories: “Old Bruise”, “New Bruise”, “Old Scar”, and “New Scar”.
Bruises typically result from blunt force or pressure, which can arise from various factors such as handling, packing, impacts, or weight. Scars, on the other hand, are caused by sharp contact due to reasons like friction from leaves or stems, cuts from knives, crates, rough handling, laborers’ nails, or the sharp tips and crowns of neighboring bunches. Old defects are usually darker than new defects as they oxidize over time (see Figure S.1). In total, 1440 defects were annotated. (37 Old Bruise, 182 New Bruise, 387 Old Scar, 834 New Scar) While the automated detection of generic defects without classification is already an impactful use case, we chose to add this categorization as it helps supply chain managers trace back the cause of the damage in the supply chain. It is important to note that there are cases that cannot be unambiguously assigned to one of those categories, even by expert annotators. However, a sub-goal of this work is to assess to which degree an automated classification is possible. It is worth mentioning that there are plenty of other relevant surface defects reducing the quality of banana crops that are not included in this study, e.g., insect damage, speckles, sunburn, maturity stains, malformed fruits, broken necks, or chemical residues.
@@
To obtain relative defect sizes, we also annotated segmentation masks of banana fruits. We decided to annotate bananas semantically only on a pixel level, omitting instance information for two reasons: First, as images often contain multiple bunches that are partly cropped or occluded, it is challenging to differentiate instances. Second, our use case does not require us to identify instances, as we are not interested in counting single bananas or bunches. We would rather distinguish focused foreground bunches from those that may appear in the background and, therefore, annotate two types of pixel classes: “Foreground Banana” and “Background Banana”. All pixels not annotated as a banana of defect are assigned the “Background” class. We annotated banana masks using LabelStudio and Segment Anything point prompts (see Section 2.3 for more details).
2.3 Mask generation with SAM
The Segment Anything Model (SAM) is a recently published foundation model for image segmentation [15]. It was trained on a large-scale dataset comprising diverse images from various domains, including natural scenes, urban environments, medical imagery, and other contexts. It is the first model of its kind specifically tailored for generalized image segmentation and promises to facilitate the implementation of domain-specific segmentation models.
Figure 1 (left) shows a high-level overview of SAM’s components. The model consists of three major components: (1) a vision transformer [7] pre-trained with the Masked Autoencoder method [9] that serves as an image encoder; (2) a prompt encoder that allows the user to specify bounding boxes or points to localize the region of interest in an image; (3) a mask decoder that generates segmentation masks based on the concatenated image and prompt embeddings. This approach yields impressive results for zero-shot instance segmentation in various domains. However, compared to conventional segmentation models, it comes with two disadvantages that require further adaptations or post-processing to enable SAM for a typical image segmentation use case: Firstly, SAM requires weak labels, such as bounding boxes or point coordinates, as input prompts. Secondly, the output masks are generally uncategorized.
We explore the use of SAM in a zero-shot setting to generate dense panoptic masks from coarse annotations—specifically, bounding boxes for defects and point annotations for banana fruits—effectively serving as a semi-automated annotation tool. However, it is not immediately clear whether SAM-generated masks sufficiently align with human-annotated masks. Other than common objects in images, surface defects potentially have ambiguous boundaries that are subject to expert judgment.
Validating SAM masks
We validate SAM masks in two ways: First, we directly evaluate the alignment between SAM-generated and human-annotated masks under different conditions, including ablations on the backbone model size and the resolution of the defect areas in the image (Section 3.1). Second, we assess the performance of our panoptic segmentation model against both human-annotated and SAM-generated masks (Section 3.2). This allows us to draw conclusions about the extent of error propagation resulting from inaccurate mask generation. More specifically, we report panoptic evaluation using three different combinations of mask sources for training and validation data: (a) both training and validation use hand-annotated defect masks; (b) both training and validation use SAM-generated defect masks; and (c) training uses SAM-generated masks while validation uses hand-annotated masks. By comparing the evaluation metrics across these three variations, we can quantify the trade-off between the time-saving benefits of automated mask generation and the quality of the resulting output.
@@ 2.4 Image segmentation
In Section 2.1, we argued why we frame this use case as a hand as a panoptic segmentation problem. This approach yields a dense semantic segmentation mask for the whole image while instance identification is provided for certain specified categories only [4, 5, 13, 14]. In our study, we declare defects as “things” (countable) and bananas as “stuff” (non-countable) as we are not interested in distinguishing single banana fruits. We train a Maskformer [4] pre-trained on the ADE20k dataset [30] in a standard supervised learning workflow (Figure 1, right).
Evaluation metrics
We evaluate the final model using metrics from all three paradigms: semantic, instance, and panoptic segmentation. Semantic segmentation maps are evaluated against a groundtruth using the Intersection-over-Union (IoU) score for a single class or the mean IoU (mIoU) score for all classes in a dataset [8, 18]. For instance detections, we follow the established practice to report average precision at different $IoU$ threshold ($AP@50$, $AP@75$) [8]. $AP$ is equivalent to $AP^{IoU=.50:.05:.95}$, averaged over ten equidistant $IoU$ thresholds between $0.50$ and $0.95$ as proposed by Lin et al. [18]. For the average recall metric ($AR$), we follow the definition of Hosang et al. [10] evaluating a range of IOU thresholds. In addition, we use Panoptic Quality ($PQ$) as introduced by Kirillov et al. [14] as an integrated metric for panoptic segmentation.
@@
Further details on model training, such as hyperparameters and data augmentation, can be found in Appendix B.